Circular No. 8985

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2008 N2–N11, C/2008 O1, AND C/2008 O4 (SOHO) Further to IAUC 8984, additional near-sun presumed comets have been

found on SOHO website images — all Kreutz sungrazers except for C/2008 N4 (a Kracht-group member that was suggested by R. Kracht as being identical with C/2002 S7; cf. IAUC 7986, plus the discussion and analysis on MPEC 2008-P60) and C/2008 N10 (no known group). C/2008 N2 was very bright (mag ∼ 4.5) and elongated in C3 images, and showed a diffuse tail ∼ 0^{o}.5 long in C2 images (where it appeared headless inside ∼ 6R_{⊙} (apparent); SECCHI COR2-A images on July 3–4 showed a thin tail much like that seen in C2 images, and a tail also was seen to form in the last images obtained with the COR2-B camera on July 4. C/2008 N3 was small and slightly diffuse (mag ∼ 7.5). C/2008 N4, C/2008 N10, and C/2008 N11 were small and stellar in appearance (mag ∼ 7.5). C/2008 N5 was quite faint (mag ∼ 7.5) and stellar in appearance in C3 images, while it was slightly diffuse with no tail in C2 images. C/2008 N6 was small and stellar in appearance (mag ∼ 7.5) in C3 images, but slightly diffuse in C2 images. C/2008 N7 was small and slightly diffuse (mag ∼ 8). C/2008 N8 was quite bright (mag ∼ 6) and slightly elongated in C3 images; it was teardrop-shaped and slightly diffuse in C2 images. C/2008 N9 was small and stellar in appearance (mag ∼ 7) in C3 images, but diffuse and slightly elongated in C2 images. C/2008 O4 was stellar in appearance, peaking at mag ∼ 5.5 around Sept. 6.83 UT.

Additional astrometry of C/2008 O1 (cf. IAUC 8962) was published on MPEC 2008-P33, with the note by K. Battams that it appeared very small and stellar in appearance, peaking at mag ∼ 6.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 N2 July 2.888 6^{h}35.^{m}8 +18^{o}46′ C3/2* BZ 2008-P60
C/2008 N3  3.913 6 44.7 +21 37 C2  RK 2008-P60
C/2008 N4  4.104 6 45.6 +23 14 C2  TH 2008-P60
C/2008 N5  4.388 6 43.7 +20 36 C3/2 AK 2008-P60
C/2008 N6  5.238 6 44.5 +19 23 C3/2 HS 2008-S67
C/2008 N7 6.104 6 53.7 +21 31 C2 TH 2008-S67
C/2008 N8  7.654 6 52.5 +18 45 C3/2 BZ 2008-S68
C/2008 N9  8.321 6 59.2 +20 23 C3  BZ 2008-S68
C/2008 N10  10.579 7 14.7 +21 52 C2  RK 2008-S68
C/2008 N11  12.513 7 15.1 +20 00 C3  HS 2008-S68
C/2008 O4  25.421 8 12.5 +20 17 C2  MK 2008-S67
2008 September 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green